# Reward-Augmented Data Enhances Direct Preference Alignment of LLMs

### v3

@@ Abstract
Preference alignment in Large Language Models (LLMs) has significantly improved their ability to adhere to human instructions and intentions. However, existing direct alignment algorithms primarily focus on relative preferences and often overlook the qualitative aspects of responses, despite having access to preference data that includes reward scores from judge models during AI feedback. Striving to maximize the implicit reward gap between the chosen and the slightly inferior rejected responses can cause overfitting and unnecessary unlearning of the high-quality rejected responses. The unawareness of the reward scores also drives the LLM to indiscriminately favor the low-quality chosen responses and fail to generalize to responses with the highest rewards, which are sparse in data. To overcome these shortcomings, our study introduces reward-conditioned LLM policies that discern and learn from the entire spectrum of response quality within the dataset, helping extrapolate to more optimal regions. We propose an effective yet simple data relabeling method that conditions the preference pairs on quality scores to construct a reward-augmented dataset. This dataset is easily integrated with existing direct alignment algorithms and is applicable to any preference dataset. The experimental results across instruction-following benchmarks including AlpacaEval 2.0, MT-Bench, and Arena-Hard-Auto demonstrate that our approach consistently boosts the performance of DPO by a considerable margin across diverse models such as Zephyr, Mistral, Qwen2, Llama3.1, Gemma2, and SPPO. Additionally, on six academic benchmarks including GSM8K, GPQA, MUSR, TruthfulQA, BBH, and ARC, our method improves their average accuracy. When applying our method to on-policy data, the resulting DPO model outperforms various baselines and achieves state-of-the-art results on AlpacaEval 2.0. Through comprehensive ablation studies, we demonstrate that our method not only maximizes the utility of preference data but also mitigates the issue of unlearning, demonstrating its broad effectiveness beyond mere dataset expansion. Our code is available at https://github.com/shenao-zhang/reward-augmented-preference.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has recently seen remarkable success in aligning Large Language Models (LLMs) to follow instructions with human intentions. In this approach, AI-generated feedback serves as a stand-in for human preferences, assessing and ranking responses to prompts to construct a preference dataset. This dataset is then utilized in preference optimization algorithms to fine-tune LLMs. Among them, direct preference alignment (Rafailov et al., 2024b; Azar et al., 2023; Zhao et al., 2023; Ethayarajh et al., 2024) that bypasses the need for an explicit reward model has garnered interest for their simplicity and cost efficiency. However, these algorithms mainly concern relative preferences and often overlook the quality of responses and their gaps, leading to limitations in their effectiveness.
@@
To address these issues, we propose learning reward-conditioned policies as a straightforward fix. By optimizing the LLM to generate responses conditioning on their qualities, the model is allowed to discern and leverage patterns within responses of varied quality. As a result, learning from both chosen and rejected responses alleviates the issue of unnecessarily unlearning high-quality rejected responses; distinguishing between varying-quality chosen responses alleviates the issue of indiscriminately accepting low-quality ones. By identifying common patterns in responses of similar quality and distinguishing them from those of differing quality, the LLM becomes more adept at generalizing to more optimal responses that are sparse in data.
@@
In experiments, we first apply our method on UltraFeedback (Cui et al., 2023) and perform DPO (Rafailov et al., 2024b) on this reward-augmented preference dataset by fine-tuning on various models, including Zephyr-7B-$\beta$ (Tunstall et al., 2023b), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023a), Qwen2-7B-Instruct (Yang et al., 2024), Llama-3.1-8B-Instruct (Dubey et al., 2024), Gemma-2-9B-It (Team et al., 2024), and SPPO (Wu et al., 2024). The results show that our method consistently boosts the performance of these models as well as their DPO models by a large margin on instruction-following benchmarks such as AlpacaEval 2.0 (Dubois et al., 2024), MT-Bench (Zheng et al., 2024), and Arena-Hard-Auto (Li et al., 2024b). Our method also improves the average accuracy on a variety of academic benchmarks (GSM8K, GPQA, MUSR, TruthfulQA, BBH, and ARC). Moreover, our findings also demonstrate an improved utility of the preference data: a subsequent round of DPO using the reward-augmented data can still significantly enhance the model fine-tuned with DPO; relabeling the binarized preference dataset with the DPO implicit reward leads to further performance gains. Additional ablation studies also suggest that our method addresses the problem of unlearning and is superior not just due to the increased dataset size. When applied to on-policy data, our method enhances the DPO model, enabling it to surpass various baselines and achieve state-of-the-art performance on AlpacaEval 2.0. Our code is available at https://github.com/shenao-zhang/reward-augmented-preference.
@@ 2 Background
Consider a language model $\pi\in\Delta_{\mathcal{Y}}^{\mathcal{X}}$ that takes the prompt $x\in\mathcal{X}$ as input and outputs the response $y\in\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ are spaces of prompts and responses, respectively. Given the prompt $x\in\mathcal{X}$, a discrete probability distribution $\pi(\cdot\mid x)\in\Delta_{\mathcal{Y}}$ is generated, where $\Delta_{\mathcal{Y}}$ is the set of discrete distributions over $\mathcal{Y}$. We define the true human preference distribution as
@@
The objective for the LLM $\pi\in\Delta_{\mathcal{Y}}^{\mathcal{X}}$ is to maximize the KL-regularized expected reward. Recent works (Azar et al., 2023; Zhao et al., 2023; Tunstall et al., 2023b; Ethayarajh et al., 2024) proposed to align the LLM directly with the preference data by deriving the preference loss as a function of the LLM by the change of variables. Among them, the Direct Preference Optimization (DPO) (Rafailov et al., 2024b) loss has the following form:
$\mathcal{L}_{\text{DPO}}(\pi;\mathcal{D}_{N})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{N}}\biggl{[}\log\sigma\biggl{(}\beta\log\frac{\pi(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\biggr{)}\biggr{]},$
@@ 3.1 Limitations of Direct Alignment from Preference
Low-Quality Chosen Responses are Indiscriminately Learned.
Reward Sparsity.
@@
Table 1: High-quality rejected responses such as $y_{2}$ can be unnecessarily unlearned: $\pi^{*}(\cdot\mid x)$ deterministically generates $y_{1}$. Reward-conditioned policies learn both responses and are easier to generalize to $g=10$ with the extracted features from $g=8$ and $g=9$.
@@ 3.2 Reward-Conditioned Policies Learn from the Full Spectrum of Response Quality
A straightforward way to address the limitations of direct alignment algorithms—specifically, their inability to account for the quality of responses—is to optimize a reward-conditioned policy. In this approach, the LLM policy is trained to generate responses corresponding to different reward values, enabling it to become aware of and adapt to these reward distinctions. By doing so, the LLM not only learns the patterns associated with the preferred responses but also retains the valuable information from the rejected ones, preventing the unlearning of high-quality rejected responses. For example, in Table 2, reward-conditioned policies learn to generate both $y_{1}$ and $y_{2}$, instead of unlearning $y_{2}$. This reward-based conditioning also enhances the model’s ability to differentiate between responses of varying quality, even if both are preferred over a rejected alternative, as illustrated in Table 2. Besides, by extracting common patterns across responses with different quality levels, the LLM becomes more generalizable and is capable of generating the highest-quality responses with reward $r_{\text{max}}$ (e.g., $r_{\text{max}}=10$), which are often sparse in the training preference data.
@@ 4 Method
$\min_{\pi}\mathbb{E}_{g,x\sim\mathcal{D}_{N},y\sim\pi(\cdot\mid x,g)}\bigl{[}(g-r(x,y))^{2}\bigr{]}=\max_{\pi}\mathbb{E}_{g,x\sim\mathcal{D},y\sim\pi(\cdot\mid x,g)}\bigl{[}R(x,y,g)\bigr{]}.$ (4.1)
To optimize the RHS of Equation (4.1), we first observe that under the new goal-conditioned reward metric $r$, for each preference pair ${x^{i},y_{w}^{i},r_{w}^{i},y_{l}^{i},r_{l}^{i}}$ in $\mathcal{D}_{N}$, we have
@@
Figure 1: Construction of the reward-augmented preference dataset.
We illustrate our method in Figure 1. For each preference pair with index $i$ in $\mathcal{D}_{N}$, two goals are defined, corresponding to the reward values of the chosen response $y_{w}^{i}$ and the rejected response $y_{l}^{i}$. Specifically, under the first goal $g=r_{w}^{i}$, the relabeled rewards are $R(x,y_{w}^{i},g)=0$ and $R(x,y_{l}^{i},g)=-(r_{w}^{i}-r_{l}^{i})^{2}$. The original ranking of responses remains the same, except that the LLM is preference optimized conditioned on $g=r_{w}^{i}$. Similarly, under the second goal $g=r_{l}^{i}$, the relabeled rewards are $R(x,y_{l}^{i},g)=0$ and $R(x,y_{w}^{i},g)=-(r_{w}^{i}-r_{l}^{i})^{2}$. Thus, the chosen and rejected responses are reversed as $y_{l}^{i}$ and $y_{w}^{i}$, respectively. By generating preference pairs conditioned on the goal reward for both the chosen and rejected responses, we obtain a reward-augmented dataset of size $2N$. Finally, this new dataset can be used with any direct alignment algorithm, such as DPO.
@@
We provide the following theoretical guarantees for our method (see A.4 for a formal description).
Theorem 4.1 (Informal).
Let $J(\pi)=\mathbb{E}_{x\sim d_{0},y\sim\pi(\cdot|x,g^{\star})}\big{[}R(x,y,g^{\star})\big{]}$ be the performance measure, where $R$ denotes the ground-truth goal-conditioned reward function and $g^{\star}$ denotes the optimal goal. Under mild assumptions, the policy $\widehat{\pi}$ optimized from the reward-augmented DPO with a SFT regularizer satisfies that with probability at least $1-\delta$,
$J(\pi^{*})-J(\widehat{\pi})\!\leq\!\sqrt{\frac{1}{N}}\cdot\bigg{\{}\frac{\sqrt{6}}{4}\big{(}1+\exp(B)\big{)}^{2}\big{(}\big{(}C_{\mu_{\bar{\mathcal{D}}}}(\mathcal{R};\pi^{\star},\pi_{\text{sft}})\big{)}^{2}+1\big{)}\iota$
$\qquad\qquad\qquad+\mathbb{E}_{x\sim d_{0}}\Big{[}\mathrm{KL}\big{(}\pi^{\star}(\cdot|x,g^{\star})\|\pi_{\mathrm{ref}}(\cdot|x,g^{\star})\big{)}\Big{]}\bigg{\}},$ (4.2)
@@ 5 Related Work
Conditional LLM Fine-Tuning.
Conditioning LLMs during training has proven effective for aligning responses with specific human objectives. SteerLM (Dong et al., 2023b; Wang et al., 2023b) extends SFT by conditioning the LLM on the multi-dimensional annotated attributes in data, such as humor and toxicity, in order to steer model responses with user customizability. Directional Preference Alignment (DPA) (Wang et al., 2024a) proposed a variant of rejection sampling fine-tuning (Yuan et al., 2023; Dong et al., 2023a) that conditions on the direction of the multi-objective reward, i.e., a user-dependent linear combination of the reward attributes (helpfulness and verbosity in their experiments), that represents diverse preference objectives. These methods aim to train a single LLM that can flexibly adjust to various user preference profiles. On the contrary, our method targets the limitations of direct alignment algorithms by introducing reward-augmented relabeling. This also differs from Conditioned-RLFT (Wang et al., 2023a), which leverages the data source information by learning a class-conditioned policy with RL-free supervised learning. Reward-aware Preference Optimization (RPO), introduced in Nemotron-4 (Adler et al., 2024), attempts to approximate the reward gap using the implicit reward and is motivated to resolve the unlearning issues of DPO, which our work also addresses. However, we show that more limitations beyond unlearning can be simply fixed with reward-conditioned LLMs and propose an easy-to-implement data relabeling method that integrates seamlessly with any direct alignment algorithm.
6 Experiments
6.1 Reward-Augmented Data Boosts DPO Performance
We begin by conducting experiments to demonstrate that applying the proposed method on DPO to fixed offline preference datasets leads to consistent performance improvements.
Setup.
We adopt the UltraFeedback (Cui et al., 2023) preference dataset containing reward values scored by GPT-4 (LLM-as-Judge) that is ranged between $1$ and $10$ for each of the preference pairs. Our method constructs reward-augmented data by conditioning on these judge values. We fine-tune on various open-weight LLMs, including Mistral-7B-Instruct-v0.3, Qwen2-7B-Instruct, Llama-3.1-8B-Instruct, Gemma-2-9B-It, and SPPO (fine-tuned from Gemma2-9B-It). We use the DPO implementation in the Huggingface Alignment Handbook (Tunstall et al., 2023a). The hyperparameters and prompts that we use are listed in Appendix B.1.
Results.
We first report the performance of the trained models on instruction-following benchmarks that use LLM as a judge, including AlpacaEval 2.0 (Dubois et al., 2024), MT-Bench (Zheng et al., 2024), and Arena-Hard-Auto (Li et al., 2024b). The results are shown in Figure 2.
@@ Results.
Figure 2: Performance of the base models, the models trained with DPO on UltraFeedback, and the models trained with DPO on reward-augmented ultrafeedback on AlpacaEval 2.0, MT-Bench, and Arena-Hard-Auto benchmarks. The complete table is deferred to Appendix B.2.
Besides, we also evaluate the models on academic multi-choice QA benchmarks, including GSM8K (Cobbe et al., 2021), GPQA (Rein et al., 2023), MUSR (Sprague et al., 2023), TruthfulQA (Lin et al., 2021), BBH (Suzgun et al., 2022), and ARC Challenge (Clark et al., 2018). To better reflect the capabilities of LLMs, we adopt various settings for these benchmarks, including zero-shot, few-shot, and few-shot Chain-of-Thought (CoT). The results are shown in Table 3.

### v5

@@ Abstract
Preference alignment in Large Language Models (LLMs) has significantly improved their ability to adhere to human instructions and intentions. However, existing direct alignment algorithms primarily focus on relative preferences and often overlook the qualitative aspects of responses, despite having access to preference data that includes reward scores from judge models during AI feedback. Striving to maximize the implicit reward gap between the chosen and the slightly inferior rejected responses can cause overfitting and unnecessary unlearning of the high-quality rejected responses. The unawareness of the reward scores also drives the LLM to indiscriminately favor the low-quality chosen responses and fail to generalize to optimal responses that are sparse in data. To overcome these shortcomings, our study introduces reward-conditioned LLM policies that discern and learn from the entire spectrum of response quality within the dataset, helping extrapolate to more optimal regions. We propose an effective yet simple data relabeling method that conditions the preference pairs on quality scores to construct a reward-augmented dataset. The experiments across various benchmarks and diverse models demonstrate that our approach consistently boosts DPO by a considerable margin. Through comprehensive ablation studies, we demonstrate that our method not only maximizes the utility of preference data but also mitigates the issue of unlearning, demonstrating its broad effectiveness beyond mere data expansion. Our code is available at https://github.com/shenao-zhang/reward-augmented-preference.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has recently seen remarkable success in aligning Large Language Models (LLMs) to follow instructions with human intentions. In this approach, AI-generated feedback serves as a stand-in for human preferences, assessing and ranking responses to prompts to construct a preference dataset. This dataset is then utilized in preference optimization algorithms to fine-tune LLMs. Among them, direct preference alignment (Rafailov et al., 2024b; Azar et al., 2023; Zhao et al., 2023; Ethayarajh et al., 2024) that bypasses the need for an explicit reward model has garnered interest for its simplicity and cost efficiency. However, these algorithms mainly concern relative preferences and often overlook the quality of responses and their gaps, leading to limitations in their effectiveness.
@@
To address these issues, we propose learning reward-conditioned policies as a straightforward fix. By optimizing the LLM to generate responses conditioning on their qualities, the model is allowed to discern and leverage patterns within responses of varied quality. As a result, learning from both chosen and rejected responses alleviates the issue of unlearning high-quality rejected responses; distinguishing between varying-quality chosen responses alleviates the issue of indiscriminately accepting low-quality ones. By identifying common patterns in responses of similar quality and distinguishing them from those of differing quality, the LLM becomes more adept at generalizing to more optimal responses that are sparse in data.
@@
In experiments, we first apply our method on UltraFeedback (Cui et al., 2023) and perform DPO (Rafailov et al., 2024b) on this reward-augmented preference dataset by fine-tuning on various models, including Zephyr-7B-$\beta$ (Tunstall et al., 2023b), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023a), Qwen2-7B-Instruct (Yang et al., 2024), Llama-3.1-8B-Instruct (Dubey et al., 2024), Gemma-2-9B-It (Team et al., 2024), and SPPO (Wu et al., 2024). The results show that our method consistently boosts the performance of these models as well as their DPO models by a large margin on instruction-following benchmarks such as AlpacaEval 2.0 (Dubois et al., 2024), MT-Bench (Zheng et al., 2024), and Arena-Hard-Auto (Li et al., 2024b). Our method also improves the average accuracy on a variety of academic benchmarks (GSM8K, GPQA, MUSR, TruthfulQA, BBH, and ARC). Moreover, our findings also demonstrate an improved utility of the preference data: a subsequent round of DPO using the reward-augmented data can still significantly enhance the model fine-tuned with DPO; relabeling the binarized preference dataset with the DPO implicit reward leads to further performance gains. Additional ablation studies also suggest that our method addresses the problem of unlearning and is superior not just due to the increased dataset size. When applied to on-policy data, our method enhances the DPO model, enabling it to surpass various baselines and achieve state-of-the-art performance on AlpacaEval 2.0.
@@ 2 Background
Consider an LLM $\pi\in\Delta_{\mathcal{Y}}^{\mathcal{X}}$ that takes the prompt $x\in\mathcal{X}$ as input and outputs the response $y\in\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ are spaces of prompts and responses, respectively. Given $x\in\mathcal{X}$, a discrete probability distribution $\pi(\cdot\mid x)\in\Delta_{\mathcal{Y}}$ is generated, where $\Delta_{\mathcal{Y}}$ is the set of discrete distributions over $\mathcal{Y}$. We define the true human preference distribution as
@@
The objective for the LLM $\pi\in\Delta_{\mathcal{Y}}^{\mathcal{X}}$ is to maximize the KL-regularized expected reward. Recent work (Azar et al., 2023; Zhao et al., 2023; Tunstall et al., 2023b; Ethayarajh et al., 2024) proposed to align the LLM directly with the paired data by deriving the preference loss as a function of the LLM by the change of variables. Among them, the Direct Preference Optimization (DPO) (Rafailov et al., 2024b) loss has the following form:
$\mathcal{L}_{\text{DPO}}(\pi;\mathcal{D}_{N})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{N}}\biggl{[}\log$
$\qquad\qquad\sigma\biggl{(}\beta\log\frac{\pi(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\biggr{)}\biggr{]},$
@@ 3.1 Limitations of Direct Alignment from Preference
Table 1: High-quality rejected responses such as $y_{2}$ can be unnecessarily unlearned: $\pi^{*}(\cdot\mid x)$ deterministically generates $y_{1}$. Reward-conditioned policies learn both responses and are easier to generalize to $g=10$ with features extracted from $g=8$ and $9$.
Low-Quality Chosen Responses are Indiscriminately Learned.
Reward Sparsity.
@@ 3.2 Reward-Conditioned Policies Learn from the Full Spectrum of Response Quality
A straightforward way to address the limitations of direct alignment algorithms—specifically, their inability to account for the quality of responses—is to optimize a reward-conditioned policy. In this approach, the LLM policy is trained to generate responses corresponding to different reward values, enabling it to become aware of and adapt to these reward distinctions. By doing so, the LLM not only learns the patterns associated with the preferred responses but also retains valuable information from the rejected ones, preventing the unlearning of high-quality rejected responses. For example, in Table 1, reward-conditioned policies (the last two rows) learn to generate both $y_{1}$ and $y_{2}$, instead of unlearning $y_{2}$. This reward-based conditioning also enhances the model’s ability to differentiate between responses of varying quality, even if both are preferred over a rejected alternative, as illustrated in Table 2. Besides, by extracting common patterns across responses with different quality levels, the LLM becomes more generalizable and is capable of generating the highest-quality responses with reward $r_{\text{max}}$ (e.g., $10$), which are often sparse in the training data.
@@ 4 Method
Figure 1: Illustration of our method: construction of reward-augmented preference datasets followed by direct alignment algorithms.
@@
$\min_{\pi}\mathbb{E}_{g,x\sim\mathcal{D}_{N},y\sim\pi(\cdot\mid x,g)}\bigl{[}(g-r(x,y))^{2}\bigr{]}$
$\qquad\qquad=\max_{\pi}\mathbb{E}_{g,x\sim\mathcal{D},y\sim\pi(\cdot\mid x,g)}\bigl{[}R(x,y,g)\bigr{]}.$ (4.1)
To optimize the RHS of (4), we observe that under the new goal-conditioned reward metric $r$, we have
@@
We illustrate our method in Figure 1. For each preference pair with index $i$ in $\mathcal{D}_{N}$, two goals are defined, corresponding to the reward values of the chosen response $y_{w}^{i}$ and the rejected response $y_{l}^{i}$. Specifically, under the first goal $g=r_{w}^{i}$, the relabeled rewards are $R(x,y_{w}^{i},g)=0$ and $R(x,y_{l}^{i},g)=-(r_{w}^{i}-r_{l}^{i})^{2}$. The original ranking of responses remains the same, except that the LLM is preference optimized conditioned on $g=r_{w}^{i}$. Similarly, under the second goal $g=r_{l}^{i}$, the relabeled rewards are $R(x,y_{l}^{i},g)=0$ and $R(x,y_{w}^{i},g)=-(r_{w}^{i}-r_{l}^{i})^{2}$. Thus, the chosen and rejected responses are reversed as $y_{l}^{i}$ and $y_{w}^{i}$. By generating preference pairs conditioned on the goal reward for both the chosen and rejected responses, we obtain a reward-augmented dataset of size $2N$. Finally, this new dataset can be used with any direct alignment algorithm, such as DPO.
@@
Moreover, we provide the following theoretical guarantees for our method (see A.4 for a formal description).
Theorem 4.1 (Informal version).
Let $J(\pi)=\mathbb{E}_{x\sim d_{0},y\sim\pi(\cdot|x,g^{\star})}\big{[}R(x,y,g^{\star})\big{]}$ be the performance measure, where $R$ denotes the ground-truth goal-conditioned reward function and $g^{\star}$ denotes the optimal goal. Under mild assumptions, the policy $\widehat{\pi}$ optimized from the reward-augmented DPO with an SFT regularizer satisfies that with probability at least $1-\delta$,
$J(\pi^{*})-J(\widehat{\pi})\!\leq\!\sqrt{\frac{1}{N}}\cdot$
$\qquad\bigg{\{}\frac{\sqrt{6}}{4}\big{(}1+\exp(B)\big{)}^{2}\big{(}\big{(}C_{\mu_{\bar{\mathcal{D}}}}(\mathcal{R};\pi^{\star},\pi_{\text{sft}})\big{)}^{2}+1\big{)}\iota$
$\qquad+\mathbb{E}_{x\sim d_{0}}\Big{[}\mathrm{KL}\big{(}\pi^{\star}(\cdot|x,g^{\star})\|\pi_{\mathrm{ref}}(\cdot|x,g^{\star})\big{)}\Big{]}\bigg{\}},$ (4.2)
@@ 5 Related Work
Figure 2: Instruction-following performance of the base models, the models trained with DPO on UltraFeedback, and the models trained with DPO on reward-augmented UltraFeedback on AlpacaEval 2.0, MT-Bench, and Arena-Hard-Auto benchmarks. Our method demonstrates considerable improvements consistently across all benchmarks. The complete table is deferred to Appendix B.2.
Conditional LLM Fine-Tuning.
Conditioning LLMs during training has proven effective for aligning responses with specific human objectives. SteerLM (Dong et al., 2023b; Wang et al., 2023b) extends SFT by conditioning the LLM on the multi-dimensional annotated attributes in data, such as humor and toxicity, in order to steer model responses with user customizability. Directional Preference Alignment (DPA) (Wang et al., 2024a) proposed a variant of rejection sampling fine-tuning (Yuan et al., 2023; Dong et al., 2023a) that conditions on the direction of the multi-objective reward, i.e., a user-dependent linear combination of the reward attributes (helpfulness and verbosity in their experiments), that represents diverse preference objectives. These methods aim to train a single LLM that can flexibly adjust to various user preference profiles. On the contrary, our method targets the limitations of direct alignment algorithms by introducing reward-augmented relabeling. This also differs from Conditioned-RLFT (Wang et al., 2023a), which supervised learns class-conditioned policies with data source information, and (Cai et al., 2023; Chen et al., 2024) that focus on unified alignment with binarized or reward datasets. Reward-aware Preference Optimization (RPO), introduced in Nemotron-4 (Adler et al., 2024), attempts to approximate the reward gap using the implicit reward and is motivated to resolve the unlearning issues of DPO, which our work also addresses. However, we show that more limitations beyond unlearning can be simply fixed with reward-conditioned LLMs and propose an easy-to-implement data relabeling method without algorithm changes. (Pang et al., 2024) addressed DPO’s tendency to reduce the probability of the chosen response by incorporating an NLL loss. In contrast, our work focuses on a different limitation of DPO—its tendency to overlook qualitative aspects of responses—and proposes a data relabeling approach that requires no algorithm changes. It also differs from conditional sequence modeling based on SFT (Chen et al., 2021; Lloret et al., 2024). Due to the lack of textual feedback in UF, we empirically compare with the reward feedback variants of (Lloret et al., 2024), including SteerLM and DPA. Similar to our work, (Kim et al., 2024) also investigates how preference optimization can overlook qualitative aspects of responses. However, their focus is on overfitting to preference data, and they propose incorporating quality margins into the optimization objective. In contrast, our approach does not involve algorithmic modifications, but rather directly targets the limitations caused by the unawareness of the reward scores. Our work also differs from (Park et al., 2024), which introduces a constraint-based regularization term specifically aimed at mitigating verbosity bias.
6 Experiments
6.1 Reward-Augmented Data Boosts DPO Performance
We begin by conducting experiments to demonstrate that applying our method on DPO to fixed offline preference datasets leads to consistent performance improvements. Our setups are listed as follows.
Setup.
We adopt the UltraFeedback (Cui et al., 2023) preference dataset for this experiment. Specifically, UltraFeedback contains reward values scored by GPT-4 (LLM-as-Judge), which are ranged between $1$ and $10$ for each of the preference pairs. Our method constructs reward-augmented data by conditioning on these judge values. We choose to fine-tune on various open-weight LLMs, including Mistral-7B-Instruct-v0.3, Qwen2-7B-Instruct, Llama-3.1-8B-Instruct, Gemma-2-9B-It, and SPPO (fine-tuned from Gemma2-9B-It). The hyperparameters and prompts that we use are listed in Appendix B.1.
Results.
We first report the performance of the trained models on instruction-following benchmarks that use strong LLMs such as GPT-4 as judges, including AlpacaEval 2.0 (Dubois et al., 2024), MT-Bench (Zheng et al., 2024), and Arena-Hard-Auto (Li et al., 2024b). The results are shown in Figure 2.